# Media Coverage of War Victims: Journalistic Biases in Reporting on Israel and Gaza

### v2

@@ Discussion
Despite its extensive coverage, the present study has a number of limitations that should be acknowledged and that could serve as a starting point for future research. First of all, we took a conservative approach and included only highly respected and relatively “liberal” media in our analysis. This has almost certainly led to an underestimate of actual media bias.
Second, our media analysis lacks a comparison to other conflict situations in order to understand whether the biases identified here are unique to Palestine and Israel. The Ukraine war may offer such a comparison standard, given the temporal proximity. Thus, future studies should not only broaden the scope to include more conservative media (such as Fox News or Daily Mail), but also extend the analysis to other conflict situations (such as Ukraine).
Another important future development regards the extension of the analysis beyond the first year of war. It is conceivable that the pro-Israeli / anti-Palestinian bias in Western media may have declined during the second year of the ongoing genocide and that this may have contributed to the greater awareness and the shift in public opinion in the West. Investigating whether the media biases identified here did, indeed, shape public opinion, and to what degree, remains another important issue to be addressed in future research. Only longitudinal or experimental methods will be able to address this question and allow inferences about the actual impact of the biases observed here on public opinion and collective action.
@@
Data Collection
The New York Times (NYT), CNN, and Al Jazeera English (AJE) news articles were accessed via the LexisNexis Academic Web Services API (WSApi) [^1] LexisNexis NYU guide, https://guides.nyu.edu/lexisnexis-rest-api [^2] LexisNexis WSApi https://solutions.nexis.com/wsapi/, accessed through October 2024. These articles cover the period from October 7, 2023, to October 7, 2024. To ensure the inclusion of only highly relevant news articles, we excluded specialized content such as podcasts and audio transcripts. We also removed summary articles that condensed multiple topics into brief overviews. Additionally, we filtered the articles based on country and subject tag, retaining only those with content relevant to Palestine and Israel, specifically tagged with “Hamas,” “Israel,” “Gaza,” “West Bank,” and “Palestin.” (the term Palestin was used to account for the variations of the word, such as Palestinian(s) and Palestine).
@@
We estimated four separate GEE logistic regression models, one for each media outlet (AJE, BBC, CNN, and NYT) In each model, the dependent variable was the binary Instance Type, and the independent variable was Side. We specified an exchangeable working correlation structure and used robust (Huber-White) standard errors to account for within-article correlation.

### v3

@@ Discussion
Despite its extensive coverage, the present study has several limitations that should be acknowledged and could serve as a starting point for future research. First of all, we took a conservative approach and included only highly respected and relatively “liberal” media in our analysis. This has almost certainly led to an underestimate of actual media bias. Thus, future studies should broaden the scope to include more conservative media (such as Fox News or Daily Mail).
Another important future development regards the extension of the analysis beyond the first year of war. It is conceivable that the pro-Israeli / anti-Palestinian bias in Western media may have declined during the second year of the ongoing genocide and that this may have contributed to greater awareness and the shift in public opinion in the West. Investigating whether the media biases identified here did indeed shape public opinion and to what extent remains another important issue to be addressed in future research. Only longitudinal or experimental methods will be able to address this question and allow inferences about the actual impact of the biases observed here on public opinion and collective action.
@@
Data Selection and Sample Construction
Media Outlet Selection. Our analysis focuses on four major international news organizations: The New York Times (NYT), BBC, CNN, and Al Jazeera English (AJE). Outlet selection was guided by four criteria:
  • publication in English,
  • digital accessibility of archival content,
  • the presence of independent reporting infrastructure (including field correspondents), and
  • substantial global audience reach.
Audience reach and influence rankings consistently identify BBC, NYT, and CNN among the most widely consumed English-language news organizations worldwide. According to PressGazette (2025) [93], BBC, NYT, and CNN rank first, second, and fourth, respectively, in global visits among the top 50 English-language news websites. MSN, which ranks third, was excluded because it functions as a news aggregator without its own reporting staff or field reporters. Among non-Western outlets, several Indian news websites (including India Times, Hindustan Times, The Indian Express, India Today, and One India) appear high in the rankings, but their readership is overwhelmingly domestic or diaspora-based. The highest-ranked truly international non-Western outlet is AJE (46th place). This pattern is consistent with rankings produced by PRlab [94], which list CNN, BBC, and NYT as the first, second, and third most influential outlets for news coverage, while AJE—ranked sixth—is the leading non-Western organization. The prominence of these outlets is further supported by audience research such as the 2024 Digital News Report [95]. Using the share of respondents who reported seeing content from each brand in the past month as a measure of global reach, CNN (25%), BBC (22%), and NYT (12%) emerge as the top three news brands worldwide, while AJE (10%; ranked fifth) is again the highest among non-Western outlets. The audience scale is substantial: the three Western outlets alone attract an estimated average of 1.906 billion visits per month, including approximately 421 million unique visitors.
Temporal Scope. We restrict our analysis to the first year of the war (October 7, 2023 - October 6, 2024), as this period captured both the overwhelming majority of global public attention and the phase of maximal media influence. Independent indicators of information demand, including Google Trends and Wikipedia pageview statistics (Supplementary Figures 1–2), show a sharp spike in interest immediately following October 7, followed by sustained levels of attention that remained substantially elevated throughout the first year relative to the subsequent period.
This early attention surge marks the interval in which news organizations allocated exceptional resources, public reliance on media was greatest, and core interpretive frames crystallized. It is also the period containing most of the conflict’s humanitarian and military variance, when reporting was most intense and consequential for public understanding.
By contrast, coverage beyond the first year occurs under conditions of markedly lower public attention and reduced reporting intensity. As audience demand stabilizes at lower levels, coverage becomes more routinized and episodic, and editorial investment contracts. Because media effects and framing dynamics are strongest during periods of high salience and interpretive uncertainty, the first year provides the most analytically informative window for assessing systematic reporting asymmetries. Examining this interval therefore allows us to evaluate patterns of representation under conditions of maximal scrutiny, when media influence and agenda-setting capacity are at their peak.
Data Retrieval and Filtering. The New York Times (NYT), CNN, and Al Jazeera English (AJE) news articles were accessed via the LexisNexis Academic Web Services API (WSApi) [^1] LexisNexis NYU guide, https://guides.nyu.edu/lexisnexis-rest-api [^2] LexisNexis WSApi https://solutions.nexis.com/wsapi/, accessed through October 2024. These articles cover the period from October 7, 2023, to October 7, 2024. To ensure the inclusion of only highly relevant news articles, we excluded specialized content such as podcasts and audio transcripts. We also removed summary articles that condensed multiple topics into brief overviews. Additionally, we filtered the articles based on country and subject tag, retaining only those with content relevant to Palestine and Israel, specifically tagged with “Hamas,” “Israel,” “Gaza,” “West Bank,” and “Palestin.” (the term Palestin was used to account for the variations of the word, such as Palestinian(s) and Palestine).
@@
We estimated four separate GEE logistic regression models, one for each media outlet (AJE, BBC, CNN, and NYT). In each model, the dependent variable was the binary Instance Type, and the independent variable was Side. We specified an exchangeable working correlation structure and used robust (Huber-White) standard errors to account for within-article correlation.